# for Autonomous Vehicles in Simulation

### v2

@@ Abstract
Index Terms:
Scenarios generation, graph generation, dynamic temporal scene graphs, heterogeneous graph learning
@@ I Introduction
Addressing the challenges of plausibility and diversity in synthetic datasets, our method generates safety-critical traffic scenarios on-demand from user inputs, leveraging real-world data. The user simply requests a scenario and our model generates dynamic temporal scene graphs corresponding to the desired scenario, which are then being rendered in CARLA Simulation [8], as seen in fig. 1. In particular, the user defines the dynamic agents, the AV action – such as move, turn left/right, or overtake – and the scenario’s criticality level – i.e. visible, near, or near collision. Then, the relations between the agents – excluding the AV, namely ego-vehicle – are sampled from our database consisting of a collection of interactions extracted from the ROad event Awareness Dataset (ROAD) [9]. These serve as seeds for our graph generation model, which learns to generate temporal links corresponding to relations between the AV and the agents and static structures. This process is guided and constrained by an ontology to ensure the plausibility of interactions and scene evolution. Our method provides a structured approach in generating complex traffic situations for evaluating AV navigation performance in diverse and complex scenarios.
@@ II Related Work
Our method tackles traffic scenario generation as a graph generation task through temporal graph learning. This section reviews related work in these three key areas, compared to our work.
Traffic scenario generation. Scenario generation is a key component of applications such as extracting scenarios from naturalistic driving data or adjusting scenario parameters [10]. However, existing methods are often limited in applicability and lack plausibility. Chat-Scene [7] employs a Large Language Model (LLM) to generate safety-critical scenarios in simulation but is restricted by user prompts and lacks real-world grounding. CC-SGG [11] perturbs scene graphs to generate corner cases but relies on a small, simulation-based dataset and doesn’t capture temporal patterns. TrafficGen [6] augments scenarios on HD maps and loads them in simulation. However, it relies on artificially generated agents, reducing scenario plausibility. Similarly, real-to-synthetic approaches [12] represent scenes as graphs and generate actors’ positions on HD maps to construct traffic scenarios, but do not directly render them in simulation. Rule-based methods [13, 14] generate scenarios via ontologies but often lack relevance to real traffic events. Our method overcomes these limitations by leveraging real-world data to capture complex interactions between scene participants and generate ontology-constrained temporal scene graphs, which are then rendered in simulation.
@@ III-A Notation and Problem Definition
We define a traffic scenario as $\mathbb{S}:\{S_{t}\;|\;t\in[1,N]\}$ consisting of $N$ sequential scenes $S_{t}$. Each $S_{t}$ is represented as a scene graph $G_{t}=\langle V_{t},E_{t}\rangle$, where nodes $V_{t}$ represent key dynamic and static entities (e.g. cars, pedestrians, or lanes) and the edges $E_{t}$ denote their relationships (e.g. distance or topology). Each node $n^{i}_{t}\in V_{t}$ corresponds to a semantic category (e.g., ego-vehicle, road, pedestrian, or cyclist) with a node feature representation $h_{t}^{i}$. Similarly, each edge $e_{t}^{ij}\in E_{t}$ represents a semantic relationship (e.g., distance, topology) between two nodes $n^{i}_{t},n_{t}^{j}\in V_{t}$ with an associate edge feature representation $p_{t}^{ij}$. Edges can connect two entities or link an entity to itself, forming self-edges. This creates a collection of $N$ graphs, $\mathbb{G}:\{G_{t}\;|\;t\in[1,N]\}$ for each scenario $\mathbb{S}$. In our proposed method, the goal is to learn to generate scenario scene graphs, $\mathbb{G}$, from seed scene graphs $\tilde{\mathbb{G}}:\{\tilde{G_{t}}\;|\;t\in[1,N]\}\text{, where each }\tilde{G_{t}}\subsetneq G_{t}$. The nodes $V_{t}$ and predicted relationships in $E_{t}$ from the generated scene graphs are used to create scenarios in simulation.
@@ IV-B Temporal Scene Graph Generation
Each input graph entity is encoded to generate compact feature vector representations. We apply a set of Multi-Layer Perceptrons (MLPs) to both the node and edge features of the augmented scene graphs to learn their corresponding embedding representations. In particular, node features are encoded as $z_{n}=MLP_{n}(h)$, actors-structures edge features as $z_{a}=MLP_{e}(p_{a})$, and the ego-vehicle-related edge features as $z_{e}=MLP_{e}(p_{e})$. We then apply two layers of the Graph Attention Network (GAT) to aggregate the feature vectors of neighbouring nodes and corresponding edge features in $z_{a}$ to generate query node embeddings, $z_{n}^{\prime}=GAT(z_{n},z_{a},e_{a})$.
@@ V-A Dataset
For this work, we used the ROAD dataset [9], an extension of the Oxford RobotCar Dataset [27], to train and evaluate our approach, as well as build our agents-structures database. This dataset is designed to test an AV ’s ability to detect various road events. Video sequences were downsampled 5-fold from 12 frames per second (fps) to increase variation across 5-frame sequences, similar to [23]. Each sequence corresponds to a specific action of the ego-vehicle, defined as AV Action in table I (4). We extract the proximity relations between actors and the ego-vehicle by computing and thresholding the box-wise median Relative Inverse Depth (RID) obtained from an off-the-shelf Midas (DPT Large) network [28], with Near Collision ($<5$ $m$), Near ($5-10$ $m$), and Visible ($>10$ $m$). We generate $6,771$ temporal graphs grouped by action, where each action group is split into training, validation, and testing sets using a 70-20-10 allocation for balanced evaluation.
@@ V-B Training
To ensure stable learning and effective generalisation, we carefully tuned hyperparameters and employed several training techniques. For hyperparameters tuning, we used k-fold cross-validation with $k=5$. The dimensions of the MLP layers in our model were configured as follows: $MLP_{n}:[22,8,1]$, $MLP_{e}:[27,8,1]$, $MLP_{tr}:[3,16,1]$, and $MLP_{gat}=[64,128,256]$. The first GAT layer used dimensions $[1,64]$, followed by another GAT layer $[256,1]$. Finally, the GCN was set with dimensions $[1,1]$ to capture recurrent node aggregations. This layered structure allowed to process efficiently both node and edge features and generate the required triplet embedding representations. We additionally used a sinusoidal positional encoding of length $2$ and the Adam optimiser with $0.01$ learning rate. Gradient clipping at ±$\pm 1$ prevented exploding gradients, while Xavier initialisation helped stabilise weight distributions across layers. We also employed weight decay with $\lambda=10^{-5}$ to mitigate overfitting.
@@ VI-C Scenario Generation
[…]
@@
[…]
@@
Figure 4: Qualitative evaluation of the generated scenarios in CARLA using sampled seed scene graphs from the ROAD dataset [9] for various actions and criticalities. As the dataset was recorded in a left-hand drive country, we adjusted the topological information to align with the configuration used in CARLA, a right-hand drive environment.
@@ VII Ablation
Table VIII: Prediction performance without positional encodings.
@@ VIII Conclusion
Our proposed methodology effectively generates diverse scenarios in simulation by modeling complex semantic interactions through temporal scene graphs. We evaluated the accuracy and contextual relevance of our model’s predictions, achieving superior performance to the baseline methods. Additionally, we assessed the performance of $3$ navigation agents in our generated scenarios, demonstrating their suitability as testing environments. Our results further indicate that our approach can generate scenarios on-demand, achieving the desired criticality levels for comprehensive evaluation of AVs in various diverse situations. Overall, our method establishes a robust framework for dynamic scenario generation, leveraging temporal graphs and GNNs to enhance AV testing and evaluation.
@@ Acknowledgements
We thank Valentina Muşat for reviewing the paper, and Harry Way and Pawit Kochakarn for their work on the early stages of the scene graph extraction pipelines.

### v3

# GraphSCENE: On-Demand Critical Scenario Generation for Autonomous Vehicles in Simulation

@@ I Introduction
Addressing the challenges of plausibility and diversity in synthetic datasets, our method generates safety-critical traffic scenarios on-demand from user inputs, leveraging real-world data. The user simply requests a scenario and our model generates dynamic temporal scene graphs corresponding to the desired scenario, which are then rendered in CARLA Simulation [8], as seen in fig. 1. In particular, the user defines the dynamic agents, the AV action (move, turn left/right, or overtake) and the scenario’s criticality level (visible, near, or near collision). Then, the relations between the agents – excluding the AV, namely ego-vehicle – are sampled from our database consisting of a collection of interactions extracted from the ROad event Awareness Dataset (ROAD) [9]. These serve as seeds for our graph generation model, which learns to generate temporal links corresponding to relations between the AV and the agents and static structures. This process is guided and constrained by an ontology to ensure the plausibility of interactions and scene evolution. Our method provides a structured approach in generating complex traffic situations for evaluating AV navigation performance in diverse and complex scenarios.
@@ II Related Work
Our method tackles traffic scenario generation as a graph generation task through temporal graph learning. This section reviews related work in these three key areas.
Traffic scenario generation. Scenario generation is a key component of applications such as extracting scenarios from naturalistic driving data or adjusting scenario parameters [10]. However, existing methods are often limited in applicability and lack plausibility. ChatScene [7] employs a Large Language Model (LLM) to generate safety-critical scenarios in simulation, but is restricted by user prompts and lacks real-world grounding. CC-SGG [11] perturbs scene graphs to generate corner cases but relies on a small, simulation-based dataset and doesn’t capture temporal patterns. TrafficGen [6] augments scenarios on HD maps and loads them in simulation. However, it relies on artificially generated agents, reducing scenario plausibility. Similarly, real-to-synthetic approaches [12] represent scenes as graphs and generate actors’ positions on HD maps to construct traffic scenarios, but do not directly render them in simulation. Rule-based methods [13, 14] generate scenarios via ontologies but often lack relevance to real traffic events. Our method overcomes these limitations by leveraging real-world data to capture complex interactions between scene participants and generate ontology-constrained temporal scene graphs, which are then rendered in simulation.
@@ III-A Notation and Problem Definition
[…]
@@ IV-B Temporal Scene Graph Generation
Each input graph entity is encoded to generate compact feature vector representations. We apply a set of Multi-Layer Perceptrons (MLPs) to both the node and edge features of the augmented scene graphs to learn their corresponding embedding representations. In particular, node features are encoded as $z_{n}=MLP_{n}(h)$, actors-structures edge features as $z_{a}=MLP_{e}(p_{a})$, and the ego-vehicle-related edge features as $z_{e}=MLP_{e}(p_{e})$. We then apply two layers of the Graph Attention Network (GAT) [27] to aggregate the feature vectors of neighbouring nodes and corresponding edge features in $z_{a}$ to generate query node embeddings, $z_{n}^{\prime}=GAT(z_{n},z_{a},e_{a})$.
@@ V-A Dataset
For this work, we used the ROAD dataset [9], an extension of the Oxford RobotCar Dataset [28], to train and evaluate our approach, as well as build our agents-structures database. This dataset is designed to test an AV ’s ability to detect various road events. Video sequences were downsampled 5-fold from 12 frames per second (fps) to increase variation across 5-frame sequences, similar to [23]. Each sequence corresponds to a specific action of the ego-vehicle, defined as AV Action in table I (4). We extract the proximity relations between actors and the ego-vehicle by computing and thresholding the box-wise median Relative Inverse Depth (RID) obtained from an off-the-shelf Midas (DPT Large) network [29], with Near Collision (<5<5 mm), Near (5105-10 mm), and Visible (>10>10 mm). We generate $6,771$ temporal graphs grouped by action, where each action group is split into training, validation, and testing sets using a 70-20-10 allocation for balanced evaluation.
@@ V-B Training
To ensure stable learning and effective generalisation, we carefully tuned hyperparameters and employed several training techniques. For hyperparameters tuning, we used k-fold cross-validation with $k=5$. The dimensions of the MLP layers in our model were configured as follows: $MLP_{n}:[22,8,1]$, $MLP_{e}:[27,8,1]$, $MLP_{tr}:[3,16,1]$, and $MLP_{gat}=[64,128,256]$. The first GAT layer used dimensions $[1,64]$, followed by another GAT layer $[256,1]$. Finally, the GCN was set with dimensions $[1,1]$ to capture recurrent node aggregations. This layered structure allowed to process efficiently both node and edge features and generate the required triplet embedding representations. We additionally used a sinusoidal positional encoding of length $2$, the Adam optimiser with $0.01$ learning rate, and weight decay with $\lambda=10^{-5}$. Gradient clipping at $\pm 1$ prevented exploding gradients, while Xavier initialisation helped stabilise weight distributions across layers.
@@ VI-C Scenario Generation
To evaluate the suitability of our generated simulation scenarios for AV testing, we assess the performance of $3$ CARLA agents – normal, cautious, and aggressive – provided by the simulator library. Each agent is tested in $30$ randomly selected scenarios with evenly distributed criticality levels. We compute the Scenario Consistency Rate (SCR) $(\%)$, measuring whether the agents’ final states align with each scenario’s criticality, as shown in table V. Our results demonstrate that our method effectively generates diverse scenarios. The behaviour of the CARLA agents in those scenarios mostly leads to situations matching the intended criticality, especially for visible and near scenarios. Interestingly, the behaviour of the aggressive agent consistently led to more dangerous situations than expected, even in lower-criticality scenarios like visible and near. We also note a lower SCR percentage for near collision scenarios, likely due to the fact that the agents in CARLA have access to the perfect traffic state – i.e. knowing the state of all vehicles and pedestrians – making it easier to avoid collisions.
@@
Table V: Performance evaluation of $3$ CARLA agents, measured in SCR $(\%)$ for assessing alignment with scenario criticality. Predicted scene graphs define the behaviours and positions of non-ego-vehicle entities, as the ego-vehicle navigates the generated scenario.
@@
Figure 4: Qualitative evaluation of the generated scenarios in CARLA using sampled seed scene graphs from the ROAD dataset [9] for various actions and criticalities. As the dataset was recorded in a left-hand drive country, we mirrored the topological information to align with the configuration used in CARLA, a right-hand drive environment.
@@ VII Ablation
Table VIII: Performance without positional encodings.
@@ VIII Conclusion
Our proposed methodology effectively generates diverse scenarios in simulation by modelling complex semantic interactions through temporal scene graphs. We evaluated the accuracy and contextual relevance of our model’s predictions, achieving superior performance to the baseline methods. Additionally, we assessed the performance of $3$ navigation agents in our generated scenarios, demonstrating their suitability as testing environments. Our results further indicate that our approach can generate scenarios on-demand, achieving the desired criticality levels. Overall, our method establishes a robust framework for dynamic scenario generation, leveraging temporal graphs and GNNs to enhance AV testing and evaluation in diverse scenarios.
@@ Acknowledgements
We thank Valentina Muşat for reviewing the paper prior to submission, and Harry Way and Pawit Kochakarn for their work on the early stages of the scene graph pipelines.